Reduce the equation into canonical form and find the solution

sin2x uxx+2cosx uxy - uyy=0


1
Expert's answer
2021-01-15T13:47:48-0500

"sin^2x.u_{xx}+2cosx.u_{xy}-u_{yy}=0\;\;\;\;...(1)\\
A=sin^2x,\;B=2cosx,\;C=-1"

"B^2-4AC=4(cos^2x+sin^2x)=4>0"

So, (1) hyperbolic.


"A\lambda^2-B\lambda+C=0\\
\Rightarrow \lambda=cosx+1\;\;or, \;cosx-1\\\;\\
\therefore \frac{dy}{dx}=cosx+1\;or, \;cosx-1\\
\Rightarrow y=sinx+x+c_1\;\;or, \;sinx-x+c_2\\
let, \phi(x,y)=y-sinx-x ,\;\;\varphi(x,y)=y-sinx+x"


Let, "\zeta=\phi(x,y), \;\; \eta=\varphi (x,y)"

So,

"\zeta =\phi(x,y)=y-sinx-x\\
\zeta_x=-cosx-1\\
\zeta _y=1\\
\zeta_{yy}=0\\
\zeta _{xx}=sinx \\
and,\\
\eta =\varphi(x,y)=y-sin x+x\\
\eta _x=-cosx+1\\
\eta_y=1\\
\eta_{yy}=0\\
\eta_{xx}=sinx"


Now,

"u_{xx} = u_{\zeta}(\zeta_x^2) + 2u_{\zetaη}\zeta_xη_x + u_{ηη}η_x^2 + u_{\zeta}\zeta_{xx} + u_ηη_{xx}\\
u_{xy} = u_{\zeta\zeta}\zeta_x\zeta_y + u_{\zetaη}(\zeta_xη_y + \zeta_yη_x) + u_{ηη}η_xη_y + u_\zeta\zeta_{xy} + u_ηη_{xy}\\
u_{yy} = u_{\zeta\zeta}\zeta_y^2 + 2u_{\zetaη}\zeta_yη_y + u_{ηη}η_y^2 + u_\zeta\zeta_{yy} + u_ηη_{yy}"


Putting above values in "(1)",

we get,

"\left(cos^4x+2cos^2x+3\right)u_{\zeta\eta}\\
+u_{\zeta\zeta}(cos^4x+2cos^3x+2cos^2x)\\+u_{\eta\eta}(cos^4x-2cos^3x+2cos^2x)=u_{\zeta}sin^3x\;\;\;...(2)\\"


From expressions of "\zeta \;and\;\eta,"

"x=\frac{1}{2}(\eta-\zeta)\\"

Now from (2),

"\left(cos^4\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta)+3\right)u_{\zeta\eta}\\
+u_{\zeta\zeta}(cos^4\frac{1}{2}(\eta-\zeta)+2cos^3\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta))\\+u_{\eta\eta}(cos^4\frac{1}{2}(\eta-\zeta)-2cos^3\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta))\\=u_{\zeta}sin^3\frac{1}{2}(\eta-\zeta)\;\;\;...(2)\\"


Hence the required canonical form is :




"\left(cos^4\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta)+3\right)u_{\zeta\eta}\\
+u_{\zeta\zeta}(cos^4\frac{1}{2}(\eta-\zeta)+2cos^3\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta))\\+u_{\eta\eta}(cos^4\frac{1}{2}(\eta-\zeta)-2cos^3\frac{1}{2}(\eta-\zeta)+2cos^2\frac{1}{2}(\eta-\zeta))\\=u_{\zeta}sin^3\frac{1}{2}(\eta-\zeta)\;\;\;"